Dan and Alice noticed that each flock passing overhead was led by one bird. In the next row were two birds, then three birds, then four birds, and so on. Alice counted forty-five birds in one flock. how many rows were there? What is the algrbra equation to find the solution?
Answer by stanbon(75887)

Dan and Alice noticed that each flock passing overhead was led by one bird. In the next row were two birds, then three birds, then four birds, and so on. Alice counted forty-five birds in one flock. how many rows were there? What is the algebra equation to find the solution?
----------------------
Equations:
S(n) = (n/2)(a + L)
---
Comment: n and L must be whole numbers.
If you let L = 9 you can solve for for "n".
---
45 = (n/2)(1 + 9)
---
90 = n(10)
n = 9 (number of rows)
=======================
Cheers,
Stan H.
=======================